OpenStudy (anonymous):

2(r + 3) – 2 = 5r + 16

OpenStudy (anonymous):

Add 2 to both sides and distribute. 2r + 6 = 5r + 18 Simplify. -12 = 2r -3 = r

OpenStudy (jojo4eva):

2r+6-2=5r+16 2r+4=5r+16 4=3r+16 -12=3r -4=r

OpenStudy (anonymous):

My bad, I say 16 and thought it was 18.

OpenStudy (jojo4eva):

the answer is -4 not -3

OpenStudy (anonymous):

And the fact that magically 5 - 2 = 2 XD